OpenStudy (anonymous):

Rewrite the following equation in point-slope form. 5x-2y=3

OpenStudy (anonymous):

point slop form is y-y1=m(x-x1)

OpenStudy (anonymous):

I know, I still cannot figure it out though.

OpenStudy (anonymous):

feather...

OpenStudy (anonymous):

feather...

OpenStudy (anonymous):

flutter wingspan retricewingspan sucking pussy

OpenStudy (anonymous):

well x= 3/5 and y would = -3/2. but they did not give you a set of co-ordinates